The sum of 2 numbers is 36. If 3 times the smaller number is subtracted from the larger number, the result if 4. : The sum of 2 numbers is 36. If 3 times the smaller number is subtracted from the larger number, the result if 4.
Answer by checkley71(8405)

X+Y=36 OR Y=36-X
Y-3X=4 OR Y=3X+4
36-X=3X+4
-X-3X=4-36
-4X=-32
X=-32/-4
X=8 ANSWER.
Y=36-8
Y=28 ANSWER.
PROOF:
28=3*8+4
28=24+4
28=28